Hey gang, I have 2 problems on perfect squares:
1. Find the value of "d" such that t^2 - 14t + d is a perfect square.


2. Find the value of "k" such that z^2 - 2z + k is a perfect square.


Thanks

Answer by checkley71(8403)

t~2-14t+d or t~2-14t+49 or (t-7)(t-7)
z~2-2z+k or z~2-2z+1 or (z-1)(z-1)
solution for all these type of problems is to square 1/2 of the middle term digit thus 14/2=7 & 2/2=1 and square them 7*7=49 1*1=1